OpenStudy (mathmath333):

whats the units digit for (52^97)*(43^72)

OpenStudy (ikram002p):

like how many digits the answer will be ? or what is first digit ?

OpenStudy (mathmath333):

hundred,ten,unit..........units digit is needed (one digit)

OpenStudy (ikram002p):

what lol ?

OpenStudy (anonymous):

Dude man there's like 285 decimal places

OpenStudy (ikram002p):

haha bat :P

OpenStudy (ikram002p):

ok just determine which digit you want to find hehe :P

OpenStudy (anonymous):

10^284 idk what it's called

OpenStudy (mathmath333):

lets assume ans will be .....88999812 here units digit is 2

OpenStudy (dls):

115359438714681346382145850061896213461923232053969372763963859870671636459648757732966007126670442160564203941945826884169658186442957982512790889307915775164917252041299013231302130473445843179328581228029719622838038897741862629063384847601854612992902492312967602215395516924035072 Answer is 2. :"D

OpenStudy (ikram002p):

ohhh lol i asked u that :o so u wanna first digit :P ok (52^97)*(43^72) mod 10

OpenStudy (anonymous):

xD

OpenStudy (ikram002p):

yeah this is bat :P can u solve it ?

OpenStudy (ikram002p):

DLS :P

OpenStudy (ikram002p):

brb ill solve it using mod thingy

OpenStudy (anonymous):

Tis 2

ganeshie8 (ganeshie8):

you need to take mod 10 right ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

take it

ganeshie8 (ganeshie8):

you have done several problems like this before, right ?

ganeshie8 (ganeshie8):

(52^97)*(43^72) mod 10 reduce

OpenStudy (mathmath333):

yes but it is still confusing

OpenStudy (mathmath333):

2^97 and 3^72

ganeshie8 (ganeshie8):

thats a good start :)

ganeshie8 (ganeshie8):

(52^97)*(43^72) mod 10 (2^97)*(3^72) mod 10

ganeshie8 (ganeshie8):

whats next

OpenStudy (mathmath333):

(2^19)(4)*(7^24)

OpenStudy (ikram002p):

oh the question been taken xD

OpenStudy (mathmath333):

(2^3)(2^6)*(-1)

ganeshie8 (ganeshie8):

step by step. see if below makes sense : (52^97)*(43^72) mod 10 (2^97)*(3^72) mod 10 (2^97)*(3^2)^36 mod 10 (2^97)*(-1)^36 mod 10 (2^97) mod 10

OpenStudy (anonymous):

Ganeshie for OS presidency xD

OpenStudy (ikram002p):

-.-

OpenStudy (anonymous):

<3

OpenStudy (ikram002p):

i only went for 2 mnts xD and come back found it taken lol

OpenStudy (ikram002p):

moshera mo monsheran meh damoh

OpenStudy (mathmath333):

u ingored (-1) because its even power?

ganeshie8 (ganeshie8):

(-1) was not ignored. (-1)^36 = +1

OpenStudy (mathmath333):

ohh

ganeshie8 (ganeshie8):

Oh sorry i skipped a step, see if below looks okay : (52^97)*(43^72) mod 10 (2^97)*(3^72) mod 10 (2^97)*(3^2)^36 mod 10 (2^97)*(-1)^36 mod 10 `(2^97)*(+1) mod 10` (2^97)mod 10

ganeshie8 (ganeshie8):

@ikram002p stop cussing >.< show him how to reduce (2^97)mod 10

OpenStudy (ikram002p):

u started -.- u show him

ganeshie8 (ganeshie8):

@mathmath333 happy with the reduction so far ?

OpenStudy (mathmath333):

yes its coming 2

ganeshie8 (ganeshie8):

looks good !

OpenStudy (mathmath333):

yes

OpenStudy (ikram002p):

as a reduce in case u wanna something less than 10 (2^97)mod 10=(2^97)mod 5 ( since 10=2*5) and (2^97 =0 mod 2) so work on this (2^97)mod 5 just for information :P

OpenStudy (dls):

what is all this mod mod :/

OpenStudy (mathmath333):

yes